# Cost and accuracy of long-term graph memory in distributed LLM-based multi-agent systems

### v1

@@ Abstract
Distributed multi-agent systems use large language models to enable collaborative intelligence while preserving privacy, yet systematic evaluations of long-term memory under network constraints remain limited. This study presents a flexible testbed comparing mem0, a vector-based memory framework, and Graphiti, a graph-based knowledge graph, using the LOCOMO long-context benchmark. Experiments were conducted under unconstrained and constrained network conditions, measuring computational, financial, and accuracy metrics. Results indicate that mem0 significantly outperforms Graphiti in efficiency, with faster loading times, lower resource consumption, and minimal network overhead, while accuracy differences are not statistically significant. Applying a statistical pareto efficiency framework, mem0 is identified as the optimal choice that balances cost and accuracy in DMAS.
Keywords: distributed multi-agent systems, long-term memory, large language models, benchmarking, vector database, knowledge graph
Open-source research: the source code of the testbed, experimental results as well as notebooks for analyzing and evaluating the results are available at: https://github.com/wolffbe/dmas-long-context-memory.
@@ 1 Introduction
Artificial intelligence (AI) agents based on large language model (LLM) are capable of methodically translating goals into actions, accessing tools interfacing third-party services, and learning from past experiences autonomously retrieved from memory [87].
By connecting multiple LLM-based agents, multi-agent system (MAS) achieve a higher form of collective intelligence that has been applied in various fields and industries [22, 59, 85, 62, 32, 30, 17], e.g. poetry [84], construction [34] and finance [71].
distributed multi-agent system (DMAS) emerged from a need to mitigate privacy risks as centralized parties in MAS process sensitive user information [75]. The vision is a global network of DMAS that collaborate using local private memory to complement knowledge-intensive human labor, research and innovation [3].
Long-term memory benchmarks such as LOCMO exist to evaluate the contextual learning ability of LLM [38] and MAS [68]. Various memory providers have already proven to enhance contextual learning using such benchmarks [68, 49, 39, 12]. Recent research has also been focused on identifying best practices and design patterns to increase the cost-effectiveness of MAS [73].
However, systematically evaluating the contextual learning capabilities of novel memory frameworks combined with newly researched best practices and design patterns for MAS remains challenging. Memory frameworks are often assessed solely on accuracy and token usage without integration into actual MAS, overlooking additional computational costs and latency. New design patterns are typically evaluated in domain-specific implementations that lack repeatable and comparable performance tests across different MAS and benchmarks. Finally, no research exists so far that systematically examines and enhances long-term memory in DMAS, particularly across various network scenarios.
@@ 1.1 Theoretical framework and literature study
A review of 67 papers addressing LLM-based agents, MAS and DMAS revealed five common fields of research across all three agent types, namely technology, optimization, safety and security, networking and memory. No existing literature on safety and security for LLM-based agents and memory for DMAS was identified during this review. This chapter highlights the theoretical framework and literature that provide the foundation to this, presumably, first research of memory for DMAS.
@@
Use cases for agents in networking are network slicing [60], viewport prediction, adaptive bitrate streaming, cluster job scheduling [70], optimizing networks in Open Radio Access Network environments [52], mitigating data exfiltration scenarios [51], and managing software-defined networks [5]. Another use case is the distributed inference of LLM at the edge of a network by dividing the model into partitions, which are then assigned to different devices [37], for example, mobile phones [42]. The placement and migration of such distributed LLM have also been managed using ant colony algorithms [67].
@@
MAS are categorized based on actors, types, communication structures, collaboration strategies, and coordination mechanisms [61]. Structures range from centralized to fully decentralized as well as static to dynamic implementations [74]. Similarly to agent memory benchmarks, MAS structure benchmarks such as AgentsNet exist to assess their performance [20].
@@
Safety and security is named as a key research direction in the field of MAS [58]. Failures in MAS have led to specification issues, for example, disobeying task or role specifications, inter-agent misalignment such as failing to ask for clarification, withholding information or ignoring other agent’s inputs, and task verification, e.g. premature termination or incomplete, incorrect or no verification [9]. Moreover, adversarial debates, voting mechanisms and error logs mitigate the risk of agents hallucinating [73]. Graph networks have also been employed in this context to detect and defend malicious agents and implement structural interventions [66]. In general, chain and cycle structures offer a high degree of protection against the spread of malicious information [77].
Use cases of MAS managing other networks include anomaly detection [46], simulations [50], mediating multi-party business and network goals [10] and orchestrating one or more centralized [19] or decentralized [29] networks [47]. Orchestration approaches range from deploying mixture of experts architectures [56] and multi-agent reinforcement learning [82] handling data retrieval, collaborative planning as well as evaluation and reflection [28].
Short-term memory in MAS is defined as only existent for the duration of the ongoing interaction while long-term memory consists of historical queries and responses [23]. Similarly to LLM-based agents, hierarchical MAS memory such as G-Memory [80] and graph-based frameworks such as AGENTiGraph exist [88]. In the field of MAS long-context memory, MIRIX combines core, episodic, semantic, procedural, resource memory and knowledge vault memory with a MAS that manages knowledge updates and retrieval [68]. Alternatively, Intrinsic Memory Agents construct agent-specific heterogenous structured memories that evolve with agent outputs [78].
@@
An initial implementation of a DMAS consists of a cloud-based LLM-based agent as a central planner in combination with device-specific agents and generative agents to optimize task execution efficiency and accuracy [72]. Different networking scenarios were not further considered.
@@
In regard to DMAS safety and security, four the four most critical trustworthiness challenges include free riding, susceptibility to malicious attacks, communication inefficiencies, and system instability [86]. Consequently, security measures on an agent or intra-LLM level and system- or inter-LLM level are required. Intra-LLM security consists of strong LLM identity, authentication and authorization, context-aware access control as well as stateless and ephemeral LLM management. Inter-LLM security covers proactive maintenance, blockchain and distributed management, micro-segmentation and isolation as well as intelligent system monitoring and failure management [35].
Extending existing network management use cases in the field of MAS [48], AgentRAN proposes the AI-RAN Factory that generates new agents, enabling a distributed network to continuously develop and grow its own intelligence [16].
@@ 1.2 Research questions and hypothesis
Based on the theoretical framework and literature study, two research gaps were identified. First of all, and this issue prevails across all three agent system types, no testbed exists to comparatively evaluate the computational and financial cost of long-term memory across implementation in addition to token usage and benchmarks. Secondly, no DMAS has been built and tested taking different network scenarios into account. Consequently, an initial prototype of a MAS and DMAS long-term memory testbed is proposed, and applied to answer the following research questions:
  • Which memory framework provides the best balance between knowledge retention as well as computational and financial cost in MAS?
  • How do knowledge retention as well as computational and financial cost vary when MAS operate in a hybrid cloud–edge environment?
@@ 2.1 Empirical approach
Experiments were conducted employing a factorial design with two independent variables: the memory framework and the network profile. The unconstrained experiment added no artificial latency or bandwidth restrictions between agents, allowing evaluation under ideal network conditions. The constrained experiment introduced realistic limitations, including 200 milliseconds of latency, a bandwidth limit of eight Mbit/s (1 MB/s), and 50 milliseconds of jitter, simulating a typical datacenter-to-edge connection.
The LOCOMO long-context memory benchmark, previously used in the initial evaluation of mem0 and MIRIX [68, 12], served as the source of context. Each experiment consisted of two phases, the loading and the questions and answers (Q&A) phase. During the loading phase, all 19 sessions contained in the first of ten conversations from LOCOMO were added to mem0 and Graphiti. The JavaScript object notation (JSON) structure of a session is presented in appendix A.1. During the Q&A phase, the MAS answered 199 questions provided in LOCOMO about all 19 sessions. By adding all sessions from the first conversation, and asking all questions related to the first conversation, a robust sample size with statistical relevance was ensured.
@@ 2.2 Testbed architecture
The memory container operates in its own memory-internal subnet that also hosts the qdrant Docker container with the volume qdrant-data [40], and the neo4j Docker container with the volumes neo4j-logs and neo4j-data [79]. Furthermore, mem0 that is based on qdrant and Graphiti that is based on neo4j are both defined as services in the memory agent. Based on the configuration in the repository’s .env file, the memory agent persists memories using either the mem0 or Graphiti backend. All other testbed components run in the DMAS-network.
Questions about each conversation are issued through the coordinator’s /ask endpoint. As a result, the coordinator forwards each question to the memory agent to retrieve related memories, and sends the question together with the retrieved context to the responder. The responder then returns its answer to the coordinator, which returns the final response to the API analysis Jupyter notebook that collects all questions, answers and metrics.
The coordinator relies entirely on tools that are called by the small language model (SLM) qwen2.5:3b-instruct-q4_K_M with three billion parameters that is optimized for tool calling [45]. On the other hand, the responder uses the LLM gpt-4o-mini with approximately eight billion parameters. The system and user prompts that are invoked during this process are shown in appendix A.2, and an example of retrieved memories sent to the responder presented in appendix A.3.
@@
The analysis Jupyter notebook running in Python 3.13.2 also automates setting up the testbed using a Makefile, loading conversations, querying the MAS, and storing all results in CSV files, with one CSV generated for each combination of memory framework and constraint mode.
@@ 3 Results and analysis
During the unconstrained experiment, Graphiti required 3.96% more OpenAI tokens than mem0. During the constrained experiment, Graphiti cost 32% more tokens. The following table presents the financial cost including total tokens and the corresponding cost in USD.
@@
Table 1: financial cost per experiment, phase and memory
@@
During the constrained experiment, Graphiti used 77.2% more CPU in the cloud and 80.7% more at the edge, 54.8% more RAM in the cloud and 24% more at the edge, 430.7% more disk in the cloud and 89.7% less at the edge, and 5,268.7% more network bandwidth in the cloud and 4,295% more at the edge than mem0. This second table shows the usage of computational resources across both networks, experiments and memory frameworks.
@@
Table 2: computational cost per experiment, phase and memory
Based on the table below displaying the total time for each phase of both experiments and memory frameworks, mem0 was 86.5% faster during the loading phase and 14.7% faster during the Q&A phase than Graphiti. During the constrained experiment, mem0 was 82.8% faster during the loading phase and 14.5% faster during the Q&A phase than Graphiti.
@@
Table 3: temporal cost per experiment, phase and memory
Regarding knowledge retention and efficiency, the responder’s prompt was adjusted to answer with “i don’t know (IDK)” rather than being indecisive or inventing a response. As a result, Graphiti was 3.6% more accurate than mem0 in the unconstrained experiment. During the constrained experiment, Graphiti was 2% more accurate.
Table 4 details the absolute numbers and rates for correct, wrong, and indecisive answers. Furthermore, Figure 2 displayed on the next page visualizes these distributions, highlighting that while Graphiti offers slightly improved accuracy, both systems are dominated by conservative “IDK” responses.
@@
Figure 2: accuracy per experiment and memory
@@
memory experiment accracy 95% CI
@@
Table 5: accuracy with 95% Wilson CI
@@ 4.1 Economic analysis and total cost of ownership
Consequently, the following two figures show the TCO of each experiment.
Figure 3: financial cost of computation and tokens consumed during the unconstrained experiment
Figure 4: financial cost of computation and tokens consumed during the constrained experiment
@@ 4.2 Formalizing the balance between cost and accuracy
To define the balance between computational cost and knowledge retention, standard efficiency metrics widely adopted in LLM and Information Retrieval benchmarking were considered. Specifically, the cost per successful query (CpSQ), which quantifies the economic resources required per correct response [11], and the accuracy-efficiency score (AES), which normalizes performance against resource consumption [53]. These metrics provide a single scalar value to rank systems based on their return on investment.
However, applying these deterministic ratios to our dataset introduces a methodological risk. As shown in Table 6, the difference in accuracy between Graphiti and mem0 is not statistically significant ($p>0.05$). Calculating CpSQ or AES using raw accuracy percentages (11.1% vs 7.5%) would inadvertently treat statistical noise as a valid performance gain, potentially leading to false efficiency claims.
Consequently, simple ratio-based metrics are rejected in favor of a statistical pareto efficiency framework. Let $C(S)$ denote the Total Cost of Ownership and $A(S)$ the accuracy of a system $S$. In this multi-objective optimization problem, a system $S_{A}$ (mem0) is defined as strictly dominant over $S_{B}$ (Graphiti) if:
@@
    The difference $A(S_{B})-A(S_{A})$ is not statistically significant (Statistical Equivalence).
Under this rigorous definition, the ’balance’ is not a trade-off but a dominance condition. Since the performance differential is statistically nullified, the optimization function collapses to cost minimization.

### v2

# Cost and accuracy of long-term memory in Distributed Multi-Agent Systems based on Large Language Models

@@ Abstract
Distributed multi-agent systems (DMAS) based on large language models (LLMs) enable collaborative intelligence while preserving data privacy. However, systematic evaluations of long-term memory under network constraints are limited. This study introduces a flexible testbed to compare mem0, a vector-based memory framework, and Graphiti, a graph-based knowledge graph, using the LoCoMo long-context benchmark. Experiments were conducted under unconstrained and constrained network conditions, measuring computational, financial, and accuracy metrics. Results indicate mem0 significantly outperforms Graphiti in efficiency, featuring faster loading times, lower resource consumption, and minimal network overhead. Crucially, accuracy differences were not statistically significant. Applying a statistical Pareto efficiency framework, mem0 is identified as the optimal choice, balancing cost and accuracy in DMAS.
Keywords: distributed multi-agent systems; large language models; long-term memory; mem0; Graphiti; knowledge graphs; vector databases; cost-accuracy trade-off; LLM benchmarking
Open-source research: the source code of the testbed, experimental results as well as notebooks for analyzing and evaluating the results are available at: https://github.com/wolffbe/dmas-memory.
@@ 1 Introduction
Artificial intelligence (AI) agents based on large language models (LLMs) are capable of methodically translating goals into actions, accessing tools interfacing third-party services, and learning from past experiences autonomously retrieved from memory [87].
By connecting multiple LLM-based agents, multi-agent systems (MASs) achieve a higher form of collective intelligence that has been applied in various fields and industries [22, 59, 85, 62, 32, 30, 17], e.g. poetry [84], construction [34] and finance [71].
Distributed multi-agent systems (DMASs) emerged from a need to mitigate privacy risks as centralized parties in MASs process sensitive user information [75]. The vision is a global network of DMASs that collaborate using local private memory to complement knowledge-intensive human labor, research and innovation [3].
Long-term memory benchmarks such as LoCoMo exist to evaluate the contextual learning ability of LLMs [38] and MASs [68]. Various memory providers have already proven to enhance contextual learning using such benchmarks [68, 49, 39, 12]. Recent research has also been focused on identifying best practices and agent design patterns to increase the cost-effectiveness of MASs [73].
However, systematically evaluating the contextual learning capabilities of novel memory frameworks combined with newly researched best practices and design patterns for MASs remains challenging. Memory frameworks are often assessed solely on accuracy and token usage without integration into actual MASs, overlooking additional computational costs and latency. New design patterns are typically evaluated in domain-specific implementations that lack repeatable and comparable performance tests across different MASs and benchmarks. Finally, no research exists so far that systematically examines and enhances long-term memory in DMASs, particularly across different network scenarios.
@@ 1.1 Theoretical framework and literature study
A review of 67 papers addressing LLM-based agents, MASs and DMASs revealed five common fields of research across all three agent types, namely technology, optimization, safety and security, networking and memory. No existing literature on safety and security for LLM-based agents and memory for DMASs was identified throughout the review. The following chapter highlights the theoretical framework and literature that provide the foundation to this, presumably, first research of memory for DMASs.
@@
Use cases for agents in networking are network slicing [60], viewport prediction, adaptive bitrate streaming, cluster job scheduling [70], optimizing networks in Open Radio Access Network environments [52], mitigating data exfiltration scenarios [51], and managing software-defined networks [5]. Another use case is the distributed inference of LLMs at the edge of a network by dividing the model into partitions, which are then assigned to different devices [37], for example, mobile phones [42]. The placement and migration of such distributed LLMs have also been managed using ant colony algorithms [67].
@@
MASs are categorized based on actors, types, communication structures, collaboration strategies, and coordination mechanisms [61]. Structures range from centralized to fully decentralized as well as static to dynamic implementations [74]. Similarly to agent memory benchmarks, MAS structure benchmarks such as AgentsNet exist to assess their performance [20].
@@
Safety and security is named as a key research direction in the field of MASs [58]. Failures in MASs have led to specification issues, for example, disobeying task or role specifications, inter-agent misalignment such as failing to ask for clarification, withholding information or ignoring other agent’s inputs, and task verification, e.g. premature termination or incomplete, incorrect or no verification [9]. Moreover, adversarial debates, voting mechanisms and error logs mitigate the risk of agents hallucinating [73]. Graph networks have also been employed in this context to detect and defend malicious agents and implement structural interventions [66]. In general, chain and cycle structures offer a high degree of protection against the spread of malicious information [77].
Use cases of MASs managing other networks include anomaly detection [46], simulations [50], mediating multi-party business and network goals [10] and orchestrating one or more centralized [19] or decentralized [29] networks [47]. Orchestration approaches range from deploying mixture of experts architectures [56] and multi-agent reinforcement learning [82] handling data retrieval, collaborative planning as well as evaluation and reflection [28].
Short-term memory in MASs is defined as only existent for the duration of the ongoing interaction while long-term memory consists of historical queries and responses [23]. Similarly to LLM-based agents, hierarchical MAS memory such as G-Memory [80] and graph-based frameworks such as AGENTiGraph exist [88]. In the field of MASs long-term memory, MIRIX combines core, episodic, semantic, procedural, resource memory and knowledge vault memory with a MAS that manages knowledge updates and retrieval [68]. Alternatively, Intrinsic Memory Agents construct agent-specific heterogeneous structured memories that evolve with agent outputs [78].
@@
An initial implementation of a DMAS consists of a cloud-based LLM-based agent as a central planner in combination with device-specific agents and generative agents to optimize task execution efficiency and accuracy [72]. However, different networking scenarios were not further considered.
@@
In regard to DMAS safety and security, the four most critical trustworthiness challenges include free riding, susceptibility to malicious attacks, communication inefficiencies, and system instability [86]. Consequently, security measures on an agent or intra-LLM level and system- or inter-LLM level are required. Intra-LLM security consists of strong LLM identity, authentication and authorization, context-aware access control as well as stateless and ephemeral LLM management. Inter-LLM security covers proactive maintenance, blockchain and distributed management, micro-segmentation and isolation as well as intelligent system monitoring and failure management [35].
Extending existing network management use cases in the field of MAS [48], AgentRAN proposes the AI-RAN Factory that generates new agents, enabling a distributed network to continuously develop and grow its own network of intelligent agents [16].
@@ 1.2 Research questions and hypothesis
Based on the theoretical framework and literature study, two research gaps were identified. First of all, and this issue prevails across all three agent system types, no testbed exists to comparatively evaluate the computational and financial cost of long-term memory across implementations in addition to token usage and benchmarks. Secondly, no DMAS has been built and tested taking different network scenarios into account. Consequently, an initial prototype of a MAS and DMAS long-term memory testbed is proposed, and applied to answer the following research questions:
  • Which memory framework provides the best balance between knowledge retention as well as computational and financial cost in MASs?
  • How do knowledge retention as well as computational and financial cost vary when MASs operate in a hybrid cloud–edge environment?
@@ 2.1 Empirical approach
Experiments were conducted employing a factorial design with two independent variables: the memory framework and the network profile. The unconstrained experiment added no artificial latency or bandwidth restrictions between agents, allowing evaluation under ideal network conditions. The constrained experiment introduced realistic limitations, including 200 milliseconds of latency, a bandwidth limit of eight megabits (Mbits) or one megabyte per second (MBytes/s), and 50 milliseconds of jitter, simulating a typical datacenter-to-edge connection.
The LoCoMo long-context memory benchmark, previously used in the initial evaluation of mem0 and MIRIX [68, 12], served as the source of context. Each experiment consisted of two phases, the loading and the questions and answers (Q&A) phase. During the loading phase, all 19 sessions contained in the first of ten conversations from LoCoMo were added to mem0 and Graphiti. For reference, the JavaScript object notation (JSON) structure of a session is presented in appendix A.1. During the Q&A phase, the MAS answered 199 questions provided in LoCoMo about all 19 sessions. By adding all sessions from the first conversation, and asking all questions related to the first conversation, a robust sample size with statistical relevance was ensured.
@@ 2.2 Testbed architecture
The memory container operates in its own memory-internal subnet that also hosts the qdrant Docker container with the volume qdrant-data [40], and the neo4j Docker container with the volumes neo4j-logs and neo4j-data [79]. Furthermore, mem0, based on qdrant and Graphiti, based on neo4j are both defined as services in the memory agent. Based on the configuration in the repository’s .env file, the memory agent persists memories using either the mem0 or Graphiti backend. All other testbed components run in the DMAS-network.
Questions about each conversation are issued through the coordinator’s /ask endpoint. As a result, the coordinator forwards each question to the memory agent to retrieve related memories, and sends the question together with the retrieved context to the responder. The responder then returns its answer to the coordinator, which returns the final response to the APIs analysis Jupyter notebook that collects all questions, answers and metrics.
The coordinator relies entirely on tools that are called by the small language model (SLM) qwen2.5:3b-instruct-q4_K_M with three billion parameters that is optimized for tool calling [45]. On the other hand, the responder uses the LLM gpt-4o-mini with approximately eight billion parameters. The system and user prompts that are invoked during this process are shown in appendix A.2, and an example of retrieved memories sent to the responder is presented in appendix A.3.
@@
Connecting all components, a Makefile automates the testbed setup, and executes the analysis Jupyter notebook that loads conversations, queries the MAS, and stores all results in CSV files, with one CSV generated for each combination of memory framework and experiment. A detailed overview of testbed components and their versions is located in appendix A.4.
@@ 3 Results and analysis
During the unconstrained experiment, Graphiti required 3.96% more OpenAI tokens than mem0. During the constrained experiment, Graphiti cost 32% more tokens. Table 1 presents the financial cost including total tokens and the corresponding cost in USD.
@@
Table 1: financial cost per experiment, phase, and memory
@@
During the constrained experiment, Graphiti used 77.2% more CPU in the cloud and 80.7% more at the edge, 54.8% more RAM in the cloud and 24% more at the edge, 430.7% more disk in the cloud and 89.7% less at the edge, and 5,268.7% more network bandwidth in the cloud and 4,295% more at the edge than mem0. Table 2 details the specific usage of computational resources across both network environments, broken down by memory framework, experiment phase, and cloud versus edge consumption.
@@
Table 2: computational cost per experiment, phase, and memory
Based on Table 3 below displaying the total time for each phase of both experiments and memory frameworks, mem0 was 86.5% faster during the loading phase and 14.7% faster during the Q&A phase than Graphiti. During the constrained experiment, mem0 was 82.8% faster during the loading phase and 14.5% faster during the Q&A phase than Graphiti.
@@
Table 3: temporal cost per experiment, phase, and memory
Regarding knowledge retention and efficiency, the responder’s prompt was adjusted to answer with i don’t know (IDK) rather than being indecisive or inventing a response. As a result, Graphiti was 3.6% more accurate than mem0 in the unconstrained experiment. During the constrained experiment, Graphiti was 2% more accurate.
Table 4 details the absolute numbers and rates for correct, wrong, and indecisive answers. Furthermore, Figure 2 displayed on the next page visualizes these distributions, highlighting that while Graphiti offers slightly improved accuracy, both systems are dominated by conservative IDK responses.
@@
Figure 2: response distribution (Correct, Wrong, IDK) for mem0 and Graphiti under unconstrained and constrained network conditions.
@@
memory experiment accuracy 95% CI
@@
Table 5: accuracy rates and 95% Wilson CIs
@@ 4.1 Economic analysis and total cost of ownership
Consequently, Figure 3 illustrates the TCO breakdown for both experiments, comparing unconstrained (Figure 3(a)) and constrained (Figure 3(b)) network conditions.
(a) unconstrained network.
(b) constrained network.
Figure 3: financial cost of computation and token consumption for mem0 and Graphiti. Figures compare (a) unconstrained and (b) constrained network conditions. Costs are categorized into cloud (blue) and edge (orange) expenditures; note the logarithmic scale on the y-axis.
@@ 4.2 Formalizing the balance between cost and accuracy
To define the balance between computational cost and knowledge retention, standard efficiency metrics widely adopted in LLM and information retrieval benchmarking were considered. Specifically, the cost per successful query (CpSQ), which quantifies the economic resources required per correct response [11], and the accuracy-efficiency score (AES), which normalizes performance against resource consumption [53]. These metrics provided a single scalar value to rank systems based on their return on investment.
However, applying these deterministic ratios to the dataset introduced a methodological risk. As shown in Table 6, the difference in accuracy between Graphiti and mem0 was not statistically significant ($p>0.05$). Calculating the CpSQ or AES using raw accuracy percentages (11.1% vs 7.5%) inadvertently treated statistical noise as a valid performance gain, leading to false efficiency claims. Consequently, simple ratio-based metrics were rejected in favor of a statistical pareto efficiency framework.
Let $C(S)$ denote the Total Cost of Ownership and $A(S)$ the accuracy of a system $S$. In this multi-objective optimization problem, a system $S_{A}$ (mem0) is defined as strictly dominant over $S_{B}$ (Graphiti) if:
@@
    The difference $A(S_{B})-A(S_{A})$ was not statistically significant (Statistical Equivalence).
Under this rigorous definition, the ’balance’ was not a trade-off but a dominance condition. Since the performance differential was statistically nullified, the optimization function collapsed to cost minimization.